show work please(For this problem, do your work on a separate page(s), then upload your answer page(s) here.) Given the following DE: V – 4y
Answer #1

Pog 0 Given , y - 44 +4y=8-20 Find complementary solution fc - using homogeneous equation (D-4D+ 4) = 0 - D? 48 +420 in CD-9e4x - 427 8-ded 42. Viz. - (8-27 de = ( xe? (8-28 22; dx. . Vi= 1 axed e-le xe dà -0. using By parts... Ate Know? ; Turdisuru- 4 - 5 मरे Vi= f 2x e a ajax e a de 442 मा(४-२४ वर -2 .42 RTI VA -- - 11 21/2 = 32 dx 214 - - -कर- 8 -2 २ -4 __ फ - e-4x6425general solution y = get yp y = cietea xe + (x e te - 2x -que-ą +(-40 t telja e